OpenStudy (anonymous):

I need help with this problem. Divide . Then simplify by taking roots, if possible. √2025xy/3√3

OpenStudy (saifoo.khan):

Do u mind using the draw tool to make it?

OpenStudy (anonymous):

\[\sqrt{2025xy}/3\sqrt{3}\]

OpenStudy (anonymous):

Sorry about that. I will really appreciate any help you guys can give me.

OpenStudy (saifoo.khan):

\[\frac{15 \sqrt{xy}}{\sqrt3}\]

OpenStudy (anonymous):

Can you explain me how you do it?